OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2.

OpenStudy (anonymous):

Solve the following for x:\[\frac{6-(-3)}{x-10}=-\frac{3}{2} \]

OpenStudy (anonymous):

would i divide 3 over 2?

OpenStudy (anonymous):

cross mulitply\[(6--3) 2=-3 (x-10) \]Expand out and then solve for x.

OpenStudy (anonymous):

\[2 (6--3)=-3 (x-10) \]2(9)=30-3 x 3x = 30-18 Can you solve for x from here?

OpenStudy (anonymous):

no, im super bad at math

OpenStudy (anonymous):

3x = 30-18 x = (30-18)/3 x = 12/3 x = 4

OpenStudy (anonymous):

ohhh!!!!! i get it!

OpenStudy (anonymous):

Very good. Have a good day.